# K-COMP: Retrieval-Augmented Medical Domain Question Answering With Knowledge-Injected Compressor

### v1

@@ Appendix C Additional baseline
In Section 5, a comparison of various baselines for RAG methods based on prompt compression is presented. However, prior RAG approaches that do not utilize compression were not implemented. Although the primary focus of our study is not to exhaustively explore non-compression-based methods, we introduce additional state-of-the-art (SOTA) baselines to facilitate a more comprehensive evaluation of the relative performance of K-comp. Specifically, RAPTOR (Sarthi et al., 2024) has been adopted as a baseline, and experiments have been conducted under several predefined conditions. Firstly, RAPTOR requires a summarizing LLM because it recursively summarizes to construct a tree structure. Therefore, we utilized FineTune model to perform context summarization, thereby eliminating the penalty for training. Secondly, RAPTOR has a limit on corpus size due to constraints in our computational resources required for summarizing clustered chunks of the retrieval corpus. Given the massive scale of the MedCorp (Xiong et al., 2024) corpus, we have replaced it with a corpus consisting of the top 5 documents retrieved by retriever across the entire dataset. It is contended that these experimental conditions do not disadvantage RAPTOR. Rather, we believe that restricting the corpus to the top-5 retrieved documents may offer potential advantages by improving the relevance and quality of the retrieved information.
As shown in Table 6, K-comp has demonstrated its effectiveness in QA tasks utilizing both prompt compression and non-prompt compression approaches. Moreover, as evidenced in Table 3, K-comp further underscores its cost efficiency, particularly in comparison to non-prompt compression methods. Additionally, we have conducted the GPT-4o evaluation described in Section 5, and the results are reported in Table 7. The outcomes indicate that the passages retrieved by RAPTOR approach are more frequently preferred in BioASQ than other datasets. As illustrated in Table 1, the FineTune model is optimally tuned for BioASQ compared to other datasets. This indicates that both overarching concepts and detailed information within the passages are effectively integrated during the tree construction process.
Notwithstanding, K-comp demonstrates higher accuracy than the state-of-the-art retrieval-augmented QA baseline (non-prompt compression), even though RAPTOR has an advantage over the methods presented in that it uses a corpus consisting of only the Top-5 docs. This validates our hypothesis that knowledge injection enhances QA performance. However, RAPTOR was originally implemented using closed models (GPT family), which may be weak for our work (Gemma).
@@ Appendix D Licenses
MedQuAD, MASH-QA, BioASQ are licensed under the CC-BY-4.0, Apache-2.0, CC-BY-2.5 licenses, respectively. Gemma-2b, Nomic-embed-text-v1.5, Meta-Llama-3, Mixtral-8x7b-v0.1-AWQ, GPT-4o, MedAlpaca-13b, Meditron-70B-AWQ, Bge-reranker-large are licensed under the Gemma, Apache-2.0, Llama 3 Community, Apache-2.0, OpenAI, Creative Commons, Llama 2 Community, and MIT.
@@ Appendix I GPT-4o Evaluation Results
Appendix J Examples of Compressed Context Generated by Baselines
@@
Table 18: Examples of summaries generated by all the baselines

### v2

@@ Appendix C Additional Baseline
In Section 5, we compared various baselines for RAG methods that incorporate prompt compression. However, prior RAG approaches that do not utilize compression were not implemented. Although the primary focus of our study is not to exhaustively explore non-compression-based methods, we include additional state-of-the-art baselines to facilitate a more comprehensive evaluation of how K-comp performs relative to other approaches. Specifically, we adopt RAPTOR (Sarthi et al., 2024) as a baseline and conduct experiments under several predefined conditions. Firstly, RAPTOR requires a summarizing model because it recursively summarizes to construct a tree structure. Therefore, we use FineTune model to perform context summarization, thereby eliminating the penalty for training. Second, due to computational resource limitations, we imposed a size constraint on the retrieval corpus to enable summarization. Given the massive scale of the MedCorp (Xiong et al., 2024) corpus, we replace it with a smaller corpus consisting of the top 5 documents retrieved by the retriever across the entire dataset. We argue that these experimental conditions do not put RAPTOR at a disadvantage. Rather, we believe that restricting the corpus to the top-5 retrieved documents may offer potential advantages by improving the relevance and quality of the retrieved information.
As shown in Table 6, K-comp proves effective in QA tasks, whether using prompt compression or non-prompt compression approaches. Furthermore, as evidenced in Table 3, K-comp further underscores its cost efficiency, particularly compared to non-prompt compression methods. We also performed the GPT-4o evaluation described in Section 5, with the results presented in Table 7. The results indicate that the passages retrieved by RAPTOR approach are preferred more frequently in BioASQ than other datasets. As seen in Table 1, the FineTune model is particularly well-tuned for BioASQ, suggesting that RAPTOR effectively integrates both high-level concepts and detailed information during its tree construction process.
Despite RAPTOR’s advantage of using a corpus limited to the top-5 documents, K-comp achieves higher accuracy than the state-of-the-art retrieval-augmented QA baseline (non-prompt compression). This validates our hypothesis that knowledge injection enhances QA performance. However, it is worth noting that RAPTOR was originally designed using closed models from the GPT family, which may not align as effectively with our work, particularly when using smaller models such as Gemma.
@@ Appendix D Licenses
MedQuAD, MASH-QA, and BioASQ are licensed under CC-BY-4.0, Apache-2.0, and CC-BY-2.5, respectively. The models Gemma-2b, Nomic-embed-text-v1.5, Meta-Llama-3, Mixtral-8x7b-v0.1-AWQ, GPT-4o, MedAlpaca-13b, Meditron-70B-AWQ, and Bge-reranker-large are licensed under Gemma, Apache-2.0, Llama 3 Community, Apache-2.0, OpenAI, Creative Commons, Llama 2 Community, and MIT, respectively.
@@ Appendix I GPT-4o Evaluation Results
Appendix J Examples of Baselines
@@
Table 18: Examples of summaries generated by all the baselines.